# Unit-generated orders of real quadratic fields I. Class number bounds

### v2

@@ 4.1.2. $\Delta^{+}_{n}=n^{2}-4$
The remaining case of class number one for narrow Richaud–Degert field maximal orders not covered by the results above (Case (2) of Proposition 2.5) consists of all squarefree discriminants $\Delta=4n^{2}+1$. This case is called Chowla’s conjecture, as stated in [13]. It was settled by Biró [8] in 2003, using his method. The squarefree discriminants of this form that have class number $h_{\Delta}=1$ are $n\in\{1,2,3,5,7,13\}$, which gives $\Delta\in\{5,17,37,101,197,677\}.$
4.2. Computational results for unit-generated orders having class number one
We give computational results toward 1.4, to determine all unit-generated quadratic orders having class number $1$, which we list in Table 2. The list is complete for $\Delta<10^{10}$.
We find a total of $19$ distinct unit-generated orders of class number $1$ having discriminant $\Delta<10^{10}$. (The discriminant $\Delta=5$ is listed twice, because it appears in both families $\Delta^{+}_{n}$ and $\Delta^{-}_{n}$, but counted once.) Class numbers of these orders were computed in SageMath using the method ZZ[(D+D^(1/2))/2].class_number() for the quadratic order of discriminant D.
@@ 88 $68$ 22
Table 2. Discriminants $\Delta=n^{2}+r=(f_{\Delta})^{2}\Delta_{0}$ with $r\in\{\pm 4\}$ and class number $h_{\Delta}=1$, complete for $\Delta<10^{10}$. The list is complete for all fundamental discriminants (without the restriction $\Delta<10^{10}$). It is not known to be complete for all non-fundamental discriminants.
The fundamental discriminants in Table 2 were previously compiled by Mollin and Williams in [39, Table 2.1]. A necessary condition for $h_{\Delta}=1$ is that $h_{\Delta_{0}}=1$, but note that the associated $\Delta_{0}$ may not be a unit-generated order. In Table 2, when $\Delta=68$, the associated fundamental discriminant $\Delta_{0}=17$ has $\mathcal{O}_{\Delta_{0}}$ not a unit-generated order.
Recently Caiero and Darmon [11, Main Theorem, Theorem 25, and Theorem 28] present conditional results showing the list above is complete modulo their Conjecture 7. Their Conjecture 7 asserts that the image of RM points under $J_{E}^{+}[\tau]$ defines a global point in a ring class field, together with the truth of a Gross–Zagier formula in this generality.
@@ 5. Unit-generated orders with one class per genus
We treat ideal class groups for unit-generated orders that have one class per genus. Technically we treat a more general problem, that of determining all unit-generated orders whose class groups are of exponent $2$, that is, $\operatorname{Cl}(\mathcal{O}_{\Delta})=\operatorname{Cl}(\mathcal{O}_{\Delta})[2]$.
The one class per genus question for positive discriminants $\Delta\equiv 0\,\,\left(\operatorname{mod}\,4\right)$ is an analogue of the famous problem of Gauss of one class per genus problem for positive definite (primitive) integral binary quadratic forms of negative discriminant $\Delta=-4D$.
@@
    In Disquisitiones Arithmeticae, Gauss [22, Article 304] gave an empirical list of $65$ discriminants [^1] Gauss [22] used the determinant $D=ac-b^{2}=-\frac{1}{4}\Delta$, rather than the discriminant. of positive definite binary quadratic forms (restricted $\Delta\equiv 0\,\,\left(\operatorname{mod}\,4\right)$) having one class per genus, and asked if the list were complete. This problem is unsolved. See Stark [48] for information on its status and Lemmermeyer [36] for further history of genus theory.
@@
    Work of Leonard Eugene Dickson extended the list of discriminants $\Delta=-4D$ to include $36$ discriminants $\Delta\equiv 1\,\,\left(\operatorname{mod}\,4\right)$ having one class per genus. There are $101$ discriminants in the two lists combined; see [12].
In converting the one class per genus problem for positive discriminants from binary quadratic forms to orders of quadratic number fields, as was done by Dedekind [14, 16, 15], one must treat the narrow class group $\operatorname{Cl}^{+}(\mathcal{O}_{\Delta})$, as we explained in Section 3.3.
Here we report on computations of those unit-generated orders having wide class group of exponent $2$ (that is, $\operatorname{Cl}(\mathcal{O}_{\Delta})=\operatorname{Cl}(\mathcal{O}_{\Delta})[2]$), as well as the subset of those unit-generated orders having narrow class number of exponent $2$ (that is, $\operatorname{Cl}^{+}(\mathcal{O}_{\Delta})=\operatorname{Cl}^{+}(\mathcal{O}_{\Delta})[2]$, or “one class per genus”). We find a total of $86$ unit-generated quadratic orders having $2$-torsion class group and discriminant $\Delta<10^{10}$, which we list in Table 3 and Table 4. (The discriminant $\Delta=5$ is again listed twice.) The class groups computed in these two tables are wide class groups. In the case $\Delta=n^{2}+4$ in Table 4, all wide class groups coincide with narrow class groups, because the fundamental unit has norm 1-1.
@@ 11 $5,8,13,29,53,173,293$ 11
A list of discriminants having exponent $2$ class groups for maximal orders of extended Richaud–Degert type was previously computed by Loboutin, Mollin, and Williams [37], complete under the assumption of the generalized Riemann hypothesis. We have checked that our tables match theirs where they overlap. [^2] They list such orders in their Tables 1 and 3, which exclude those with class number $1$. Their Tables 1 and 3 include maximal orders for $\Delta=n^{2}+4$, and their list matches our Table 4 in the rows labeled $f=1$. They list maximal orders by radicand $D=m^{2}+r$, which is squarefree, with $4m\equiv 0\,\,\left(\operatorname{mod}\,r\right)$. Their Table 1 uses $D=r^{2}k^{2}+k$, so $m=rk$ and $r=k$, and we take $k=1$ and $m=r^{2}k^{2}$ odd, so $\Delta=4D$. The Table $3$ list uses $D=r^{2}k^{2}+4k$, and we take $k=1$ and $\Delta=D$. Their Table $1$ includes maximal orders with $\Delta=n^{2}-4$ and $2\,|\,\Delta$ (with $D=\Delta/4$), and it matches our Table 3 for $f=1$ in these cases. (Take k=1k=-1 in their Table 1, require $r$ to be even, and set $\Delta=4D$.) All the odd discriminant cases, with $\Delta=D$, appear in their Table 3. They list $\Delta=r^{2}k^{2}+4k$ in their Table 3, and the cases $r=1$ and odd k=(n+2)k=-(n+2) correspond to $D=n^{2}-4$.
@@ 6. Concluding remarks
The families of real quadratic orders with $\Delta=\Delta^{+}_{n}=n^{2}-4$ and $\Delta=\Delta^{-}_{n}=n^{2}+4$ “behave like” imaginary quadratic fields in terms of the rate of growth of their class number. One may ask whether other class group statistics for these families has similar behavior to the imaginary quadratic case. For example, one such statistic is the average size of the $m$-torsion, for fixed $m$, $\frac{1}{X}\sum_{n\leq X}\left|\operatorname{Cl}(\mathcal{O}_{\Delta^{\pm}_{n}})[m]\right|$ as $X\to\infty$. The case $m=2$ reduces by genus theory to a statistic on the number of distinct prime divisors of $\Delta^{\pm}_{n}$, which seems approachable.
1.4, the classification of all unit-generated quadratic orders having class number one, is an analogue of the classical problem of enumerating all imaginary quadratic discriminants of class number one. The imaginary quadratic discriminant problem for maximal orders was solved by the Heegner–Stark–Baker Theorem. Stark [48] notes that the non-maximal order case for imaginary quadratic fields is also easily solved, giving the four discriminants $\{-12,-16,-27,-28\}$, using the fact that non-maximal orders can have class number one only if the maximal order also has class number one.
1.4 is solved in the real quadratic maximal order case for unit-generated orders by 1.5. It remains open in the non-maximal order case. This problem has an extra difficulty: The associated maximal order to a unit-generated order need not be unit-generated. This possibility is realized for $\Delta=68$, where the quadratic order of discriminant $\Delta_{0}=17$ is not unit-generated. It is believed there are infinitely many fundamental discriminants of class number one, so the search space potentially has infinitely many fundamental discriminants to rule out.
1.8, the classification of unit-generated quadratic orders whose class group is $2$-torsion, subsumes a real quadratic analogue of the Gauss one class per genus problem, which itself is currently unsolved. The complete classification of unit-generated orders with $2$-torsion class group seems out of reach at this time, in parallel with the idoneal number problem being unsolved.
@@ Acknowledgements
Work of the first author was supported by NSF grant DMS-2302514. Work of the second author was partially supported by NSF grant DMS-1701576. We are grateful to the developers of SageMath, which we used to compute class numbers and other properties of real quadratic orders.

### v3

@@ 4.2. Unit-generated non-maximal orders with class number one
The computation must test the unit-generated order property. For $\Delta_{k}=5^{2k+1},$ we have $h_{\Delta_{k}}=1$ for all $k\geq 0$. However only $\Delta_{0}=5$ and $\Delta_{1}=125$ are unit-generated. Here $\Delta_{2}=3125$ in Step (S5) fails to be unit generated.
We summarize the results for unit-generated quadratic orders of class number one in Table 2.
@@ 88 $68$ 22
Table 2. Quadratic order discriminants $\Delta=n^{2}+r=(f_{\Delta})^{2}\Delta_{0}$ with $r\in\{\pm 4\}$ and class number $h_{\Delta}=1$.
There are $11$ real quadratic fields having unit-generated maximal orders having class number one, compared with $9$ imaginary quadratic fields having maximal order having class number one. There are $17$ unit-generated real quadratic orders having class number one, compared with $15$ imaginary quadratic orders having class number one.
@@ 5. Unit-generated orders with one class per genus
We treat ideal class groups for unit-generated orders that have one class per genus. Technically we treat a more general problem, that of determining all unit-generated orders whose (wide) class groups are of exponent $2$, that is, $\operatorname{Cl}(\mathcal{O}_{\Delta})=\operatorname{Cl}(\mathcal{O}_{\Delta})[2]$.
5.1. Gauss one class per genus problem
A famous problem of Gauss is the one class per genus problem for positive definite (primitive) integral binary quadratic forms of negative discriminant $\Delta=-4D$.
@@ 5.1. Gauss one class per genus problem
    In Disquisitiones Arithmeticae, Gauss [22, Article 304] gave an empirical list of $65$ discriminants [^1] Gauss [22] used the determinant $D=ac-b^{2}=-\frac{1}{4}\Delta$, rather than the discriminant. of positive definite binary quadratic forms (restricted to $\Delta\equiv 0\,\,\left(\operatorname{mod}\,4\right)$) having one class per genus, and asked if the list was complete.
@@
    Work of Leonard Eugene Dickson extended the list of discriminants to include odd discriminants $\Delta\equiv 1\,\,\left(\operatorname{mod}\,4\right)$, finding $36$ discriminants $\Delta\equiv 1\,\,\left(\operatorname{mod}\,4\right)$ having one class per genus. There are $101$ discriminants in the two lists combined; see [12].
The Gauss one class per genus problem for binary forms of negative discriminant remains unsolved. See Stark [47] for information on its status and Lemmermeyer [36] for further history of genus theory.
The one class per genus problem for binary quadratic forms was carried over to ideals of orders of quadratic number fields, by Dedekind [14, 16, 15]. For the genus theory of real quadratic orders, one must use the narrow class group $\operatorname{Cl}^{+}(\mathcal{O}_{\Delta})$ rather than the class group $\operatorname{Cl}(\mathcal{O}_{\Delta})$; see Section 5.2.
One may view the one class per genus question for unit generated orders of real quadratic fields as an analogue of the Gauss one class per genus problem, either with or without the restriction to $\Delta\equiv 0\,\,\left(\operatorname{mod}\,4\right)$.
5.2. Finiteness of unit-generated orders with one class per genus
We recall facts about the genus theory of Gauss for general binary quadratic forms, described in [25, Chapter 6.5], and its relation to the narrow class group $\operatorname{Cl}^{+}(\mathcal{O}_{\Delta})$ and (wide) class group $\operatorname{Cl}(\mathcal{O}_{\Delta})$ of a quadratic order $\mathcal{O}_{\Delta}$ of the same discriminant.
  1. (1)
  2. (2)
  3. (3)
Then its order is given by [25, Theorem 6.5.2], which we state as 5.1.
Proposition 5.1.
If $\Delta$ is a (real or imaginary) quadratic discriminant, then
We upper-bound $\mu(\Delta)$ in terms of $\left|\Delta\right|$. We have $\mu(\Delta)-1\leq\omega(\Delta)$, where $\omega(m)$ counts the number of distinct prime divisors of $m$ (without multiplicity). It is well known that $\omega(m)=O\!\left(\frac{\log m}{\log\log m}\right)$ for $m\geq 2$ [42, Theorem 2.10], so we conclude, for $\left|\Delta\right|\geq 20$,
$\mu(\Delta)-1=O\!\left(\frac{\log\left|\Delta\right|}{\log\log\left|\Delta\right|}\right).$ (5.3)
The constant in the $O$-notation is effectively computable.
Proof of 1.5.
We first consider arbitrary quadratic discriminants $\Delta$ (not necessarily unit-generated) such that $\operatorname{Cl}(\mathcal{O}_{\Delta})$ is $2$-torsion. Then by 5.1 we have
$\log\left|\operatorname{Cl}(\mathcal{O}_{\Delta})\right|\leq\log\left|\operatorname{Cl}^{+}(\mathcal{O}_{\Delta})\right|=\log\left|\mathfrak{F}_{\Delta}[2]\right|=\log 2^{\mu(\Delta)-1}$ (5.4)
Using (5.3), we obtain, for $\left|\Delta\right|\geq 20$,
$\log\left|\operatorname{Cl}(\mathcal{O}_{\Delta})\right|=O\!\left(\frac{\log\left|\Delta\right|}{\log\log\left|\Delta\right|}\right).$ (5.5)
From (5.5) it follows that, for any infinite family $\mathcal{F}$ of real quadratic orders in which $\operatorname{Cl}(\mathcal{O}_{\Delta})$ is $2$-torsion, we necessarily have
For real quadratic orders such infinite families do exist since, as noted earlier, there exist infinitely many real quadratic orders having class number $1$.
5.3. Computations of unit-generated orders having (wide or narrow) class group of exponent 22
We give lists of all unit-generated quadratic orders having wide class group of exponent $2$ (that is, $\operatorname{Cl}(\mathcal{O}_{\Delta})=\operatorname{Cl}(\mathcal{O}_{\Delta})[2]$) and having discriminant $\Delta<10^{10}$. We also list the subset of those unit-generated orders having narrow class number of exponent $2$ (that is, $\operatorname{Cl}^{+}(\mathcal{O}_{\Delta})=\operatorname{Cl}^{+}(\mathcal{O}_{\Delta})[2]$, or “one class per genus”). We find a total of $86$ unit-generated quadratic orders having $2$-torsion (wide) class group and discriminant $\Delta<10^{10}$, which we list in Table 3 and Table 4. (The discriminant $\Delta=5$ again appears twice.) In the case $\Delta=n^{2}+4$ in Table 4, all wide class groups coincide with narrow class groups, because the fundamental unit has norm 1-1.
@@ 11 $5,8,13,29,53,173,293$ 11
A list of discriminants having exponent $2$ class groups for maximal orders of extended Richaud–Degert type was previously computed by Loboutin, Mollin, and Williams [37], complete under the assumption of the generalized Riemann hypothesis. We have checked that our tables match theirs where they overlap. [^2] They list such orders in their Tables 1 and 3, which exclude those with class number $1$. Their Tables 1 and 3 include maximal orders for $\Delta=n^{2}+4$, and their list matches our Table 4 in the rows labeled $f=1$. Their Table $1$ includes maximal orders with $\Delta=n^{2}-4$ and $2\,|\,\Delta$ (with $D=\Delta/4$), and it matches our Table 3 for $f=1$ in these cases. All the odd discriminant cases, with $\Delta=D$, appear in their Table 3.
Remark 5.2.
For comparison with the imaginary quadratic case, the results of Frei [21, Theoréme 4.5] show the largest (absolute) discriminant of the $65$ even discriminants is $\Delta=2^{2}\cdot 1848=7392$, and the largest of the $36$ odd discriminants is $\Delta=3315$.
@@ 6. Concluding remarks
The families of real quadratic orders with $\Delta=\Delta^{+}_{n}=n^{2}-4$ and $\Delta=\Delta^{-}_{n}=n^{2}+4$ behave like imaginary quadratic fields in terms of the rate of growth of their class number. One may ask whether other class group statistics for these families has similar behavior to the imaginary quadratic case. For example, one such statistic is the average size of the $m$-torsion, for fixed $m$, $\frac{1}{X}\sum_{n\leq X}\left|\operatorname{Cl}(\mathcal{O}_{\Delta^{\pm}_{n}})[m]\right|$ as $X\to\infty$. The case $m=2$ reduces by genus theory to a statistic on the number of distinct prime divisors of $\Delta^{\pm}_{n}$, which seems approachable.
1.7, the classification of unit-generated quadratic orders whose class group is $2$-torsion, subsumes a real quadratic analogue of the Gauss one class per genus problem, which itself is currently unsolved. The complete classification of unit-generated orders with $2$-torsion class group seems out of reach at this time, in parallel with the idoneal number problem being unsolved.
@@ Acknowledgements
Work of the first author was supported by NSF grant DMS-2302514. Work of the second author was partially supported by NSF grant DMS-1701576.
We thank Henri Darmon for suggesting the key idea behind the proof of 4.7. We are grateful to the developers of SageMath, which we used to compute class numbers and other properties of real quadratic orders.